OpenStudy (anonymous):

7x/144 yards and 5y/12 feet equal how many inches?

OpenStudy (anonymous):

depends on the values of x and y.

OpenStudy (anonymous):

Eh they are not given. But the answer choices are A) 7x/12 +5y/4 B) 7x/12+5y C) 7x/4 +5y D) 7x/4+60y E) 7x+ 5y/4

OpenStudy (anonymous):

C) ans

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

(7x/144)(12/5y)(3feet/1yard)......cancelling or dividing... 144/12=12, 12/3=4 leaves us =(7x/4)(1/5y)\ =(7x/4)divided by(5y)\ ans

OpenStudy (anonymous):

any question ventr.?

OpenStudy (anonymous):

Can you explain in more details please i still dont get it :(